# Energy-Conscious LLM Decoding: Impact of Text Generation Strategies on GPU Energy Consumption

### v1

@@ 3.5 Strategies & Hyperparameters
We employed the decoding strategies introduced in Sections 2.2 & 2.3 in our study. All adopted strategies have been implemented in the Hugging Face [42] library. We also tuned the hyperparameters of each strategy To evaluate their impact on text quality and energy consumption across our selected tasks. The range of hyperparameter values for each decoding technique was selected based on recommendations from standard practices and recent studies [4, 5]. However, in some cases, we included unconventional hyperparameter values to investigate their impact, particularly on energy consumption. For example, in the Top-K decoding strategy, we conducted experiments with $k=1$. While $k=1$ is essentially equivalent to the greedy decoding strategy, this allowed us to examine whether differences in implementation between greedy decoding and Top-K decoding result in variations in performance or energy consumption. Table 3 lists the utilized decoding strategies and their associated hyperparameters. Our goal is not to identify the optimal decoding hyperparameters for the entire benchmark suite across different models but rather to examine the relationship between performance and energy consumption in comparable settings for each decoding strategy. By focusing on a limited subset of datasets, we aim to determine which decoding strategy is more energy-efficient during LLM inference.
4 Experimental Results
@@
We conducted an extensive benchmarking of different decoding strategies across 5 text generation tasks. The batch size is set to 1 in all experiments, as our selected text generation tasks from LM-Evaluation-Harness do not allow batch size modification for these cases. For each decoding method, we experimented multiple hyperparameters, repeating each configuration 5 times to capture variations in the LLM’s inference behavior. Through preliminary testing, we found that 5 repetitions were sufficient to capture variations. During each run, we tracked both text generation quality and energy consumption. After completing each job, we filtered the energy usage data to focus solely on the active GPU processing periods. This approach allowed us to accurately quantify the GPU energy consumption and ensured that our findings are both reliable and reflective of real-world GPU usage during text generation tasks.
@@
In the following sections, we examine the trade-offs between text generation quality and GPU energy consumption for each decoding strategy and analyze how hyperparameter tuning impacts energy usage.
4.1 Text Generation Quality vs Energy Efficiency
The text generation quality and energy consumption for various decoding strategies using the Qwen2.5 7B-Instruct model are presented in Table 1. The results were obtained using the best hyperparameter configuration that delivered the highest quality for each decoding method. The corresponding energy consumption for this optimal hyperparameter was averaged across all 5 runs and reported in watt-hours. Additionally, we evaluated the trade-off between quality and energy across different methods using an Efficiency Ratio (ER = Quality/Energy). This metric measures the generative quality achieved per unit of energy consumed, making it valuable for identifying methods that offer high quality while minimizing energy usage.
@@ 4.1 Text Generation Quality vs Energy Efficiency
Table 1: Text Generation Quality and Average Energy Consumption across Decoding Methods in their best Hyperparameter Setting using Qwen2.5 7B-Insruct with batch size=1. The associated hyperparameters are listed in Table 4. The table uses color coding to highlight key metrics: blue indicates high generative quality, red signifies high energy consumption, and green represents a high efficiency ratio.
@@
    De→En BLEU     41.09 42.71 42.65 22.97 41.03 41.76     41.02 41.26 41.11 31.53 35.68 38.39
  Wh     9.54 11.45 14.58 22.72 13.57 8.33     9.68 9.68 9.64 10.28 10.36 9.76
  ER     4.30 3.73 2.92 1.01 3.02 5.01     4.23 4.26 4.26 3.06 3.44 3.93
    En→De BLEU     26.66 29.88 29.33 19.11 26.94 26.47     26.73 26.96 26.85 18.12 24.93 24.97
  Wh     13.23 18.37 17.74 22.12 15.67 12.31     13.42 13.49 13.33 14.04 13.76 13.49
  ER     2.01 1.62 1.65 0.86 1.71 2.15     1.99 1.99 2.01 1.29 1.81 1.85
    GSM8K Exact-M     0.82 0.90 0.87 0.79 0.85 0.80     0.88 0.88 0.86 0.83 0.85 0.87
  Wh     13.51 20.32 30.75 18.11 21.53 10.63     13.27 13.58 13.81 13.79 13.82 13.48
  ER     0.06 0.044 0.028 0.043 0.039 0.075     0.066 0.064 0.062 0.06 0.061 0.064
    Code2Text-Py S-BLEU     0.86 0.84 0.83 0.92 1.16 0.88     0.90 0.90 0.92 0.90 0.90 0.93
  Wh     18.77 19.91 23.85 27.05 36.78 15.06     18.98 18.82 18.67 18.91 18.92 18.82
  ER     0.045 0.042 0.034 0.034 0.031 0.058     0.047 0.047 0.049 0.047 0.047 0.049
    Code2Text-Js S-BLEU     0.73 0.74 0.69 0.72 1.18 0.76     0.78 0.80 0.78 0.75 0.75 0.75
  Wh     18.97 20.21 22.43 26.57 36.77 15.63     19.06 19.15 18.94 18.97 18.79 19.04
  ER     0.038 0.036 0.03 0.027 0.032 0.048     0.04 0.041 0.041 0.039 0.039 0.039
In both translation tasks, Beam Search and DBS achieve the highest BLEU scores—up to about 42.7 for De→En and 29.9 for En→De—but also draw considerable power, pulling down their overall efficiency ratio. In contrast, AD delivers BLEU scores near—or on par with—those top‐performing beam‐based methods while consuming far less power, thereby attaining the best efficiency ratios. Greedy decoding, along with Temperature, Top‐k, and Top‐p sampling methods, occupies a middle ground: they offer moderately high BLEU at moderate energy costs, resulting in fairly respectable efficiency.
Similarly, in GSM8K, Beam search achieves the highest exact-match accuracy (0.90) but requires 20.32 Wh of energy, resulting in a modest efficiency ratio. In contrast, AD maintains solid accuracy (0.80) at just 10.63 Wh, yielding the highest efficiency ratio among all methods. Several stochastic methods also strike a better balance than beam-based approaches, generating high-quality texts at more moderate energy levels.
For both Python and JavaScript code summarization tasks, CS achieves the highest smoothed bleu scores, exceeding 1.0 in both cases, but also consumes the highest energy usage, putting it at a disadvantage efficiency‐wise. In both cases, CS, DoLa and DBS have the highest amount of energy consumption compared to other methods. In contrast, AD delivers near‐competitive smoothed bleu scores at a fraction of the energy, leading to the highest efficiency ratios among the tested strategies. Stochastic sampling techniques fall into a middle zone, similar to translation and math problem-solving tasks, offering moderate generation quality without the extreme energy consumption of beam-based or contrastive methods in deterministic approaches.
@@
To better demonstrate how the choice of decoding strategy affects generative quality and energy consumption, we employed the Relative Standard Deviation expressed as $\text{RSD}(\%)=\frac{\sigma}{\mu}\times 100$ mathematically. RSD quantifies the variability of the data relative to its mean value. We utilize this statistical metric to provide a normalized measure of variability in text generation quality and energy consumption using various decoding methods across different tasks. Intuitively, a lower RSD indicates greater stability and consistency, while a higher value reflects greater variability. We depicted the RSD of quality and energy consumption for each task in Figure 1 (based on the results from Table 1). It is clear that the selection of decoding strategies significantly affects the variability of energy consumption. Although in some cases, such as GSM8k, we do not observe much variation in output quality when using different decoding methods, we do see significant energy variation in this case. This highlights the importance of being more energy-conscious when choosing a decoding strategy for our task. Choosing an inappropriate strategy may not greatly reduce output quality, but it can lead to a significant increase in energy consumption.
Figure 1: Relative Standard Deviation of Text Generation Quality and GPU Energy Usage for each task across all decoding strategies using Qwen2.5-7B-Instruct.
@@
Table 2: Relative Standard Deviation (RSD) and Sharpe Ratio (SR) for each Decoding Strategy. RSD is calculated for text generation quality and average energy consumption across different hyperparameter values. SR takes into account the quality and energy consumption of all hyperparameter runs within a decoding method. The table uses color coding to highlight key metrics: blue indicates high RSD in quality, red signifies high RSD in energy consumption, and green represents a high SR.
@@
  De→En BL (RSD%)  0.31 0.46 25.96 68.62 94.99 0.49    51.21 9.22 12.30 2.99 5.32 6.83
  Wh (RSD%)  13.32 11.39 20.08 32.96 53.45 2.25    35.55 2.73 2.99 1.53 2.63 2.09
  SR  $7e^{-3}$ $8e^{-3}$ $8e^{-4}$ $e^{-3}$ $3e^{-4}$ $6e^{-2}$  $e^{-3}$ $3.6e^{-2}$ $2.8e^{-2}$ $5.1e^{-2}$ $3.3e^{-2}$ $4.5e^{-2}$
  En→De BL (RSD%)  1.30 2.16 28.86 79.02 104.95 5.41    51.50 13.69 19.64 2.18 4.22 9.11
  Wh (RSD%)  11.09 7.02 16.47 18.22 44.12 4.37    22.77 1.45 1.88 0.92 0.59 1.98
  SR  $4e^{-3}$ $6e^{-3}$ $9e^{-4}$ $e^{-3}$ $2e^{-4}$ $1.3e^{-2}$  $e^{-3}$ $3.1e^{-2}$ $1.9e^{-2}$ $3.5e^{-2}$ $7.3e^{-2}$ $2.2e^{-2}$
  GSM8K EM (RSD%)  1.85 1.42 3.26 64.70 92.36 24.21    41.12 3.07 3.05 1.43 2.11 2.02
  Wh (RSD%)  30.50 20.06 0.31 17.24 19.80 8.61    16.39 0.75 0.86 1.27 2.56 1.19
  SR  $3e^{-5}$ $4.5e^{-5}$ $2.3e^{-3}$ $3.2e^{-5}$ $8e^{-6}$ $2.04e^{-4}$  $7.7e^{-5}$ $1.9e^{-3}$ $1.7e^{-3}$ $1.1e^{-3}$ $6.3e^{-4}$ $1.3e^{-3}$
  Code2Text-Py SBL (RSD%)  4.67 6.99 0.48 6.77 12.35 1.44    11.03 3.18 3.30 3.16 2.66 4.88
  Wh (RSD%)  12.13 10.07 0.35 0.76 1.42 1.69    0.57 0.44 0.33 0.53 0.32 0.27
  SR  $8e^{-5}$ $8.4e^{-5}$ $1.7e^{-3}$ $1.2e^{-3}$ $5.2e^{-4}$ $9.1e^{-4}$  $1.9e^{-3}$ $2.5e^{-3}$ $3.3e^{-3}$ $2.2e^{-3}$ $3e^{-3}$ $3.6e^{-3}$
  Code2Text-Js SBL (RSD%)  3.63 3.27 2.12 1.26 21.19 1.08    7.20 4.57 3.44 2.39 2.70 1.53
  Wh (RSD%)  13.67 10.59 0.17 0.37 1.64 1.33    0.38 0.49 0.57 0.14 0.56 0.20
  SR  $6e^{-5}$ $6.7e^{-5}$ $2.6e^{-3}$ $2e^{-3}$ $4e^{-4}$ $9.7e^{-4}$  $2.3e^{-3}$ $1.9e^{-3}$ $1.7e^{-3}$ $3.9e^{-3}$ $1.7e^{-3}$ $3.1e^{-3}$
4.1.1 Key Findings:

### v2

@@ 3.3.4 Open-ended Generation
We employed the decoding strategies introduced in Sections 2 and 2 in our study. All adopted strategies have been implemented in the Hugging Face [54] library (v4.51.3). We also tuned the hyperparameters of each strategy to evaluate their impact on text quality and energy consumption across our selected tasks. The range of hyperparameter values for each decoding technique was selected based on recommendations from standard practices and recent studies [19, 20]. However, in some cases, we included unconventional hyperparameter values to investigate their impact, particularly on energy consumption. Table 1 lists the utilized decoding strategies and their selected hyperparameters. Our goal is to report the optimal decoding hyperparameters for our benchmark suite across different models and to examine the relationship between performance and energy consumption in comparable settings for each decoding strategy. By focusing on a limited subset of datasets, we aim to determine which decoding strategy is more energy-efficient during LLM inference.
3.4 Benchmarking Setup
@@
We conducted an extensive benchmarking of different decoding strategies across five text generation tasks. The batch size is set to 1 in all experiments to prioritize comparability. For each decoding method, we experimented multiple hyperparameters, repeating each parameter setting 5 times to ensure reliability and to capture energy variations in the LLM’s inference. During each run, we evaluated text generation performance and recorded GPU energy consumption, GPU utilization and inference latency. Through preliminary testing, we found that 5 repetitions were sufficient to capture potential variations or spikiness in GPU utilization. We then reported the average energy consumption across all runs. We also ensured that the SM and memory clock frequencies of both GPUs remained consistent across all experiments, since variations in clock frequency can directly affect GPU power draw and, in turn, our energy measurements. After completing each job, We utilized a post-processing pipeline that identifies the active GPU accelerator periods. In other words, we filtered the recorded logs to focus solely on the active GPU periods. This approach allowed us to accurately quantify the GPU energy consumption and ensured that our experiments maintained consistent and stable GPU utilization throughout the generation process without any GPU idle periods.
4 Experimental Results & Discussions
@@
We examine the trade-offs between text generation quality and GPU energy consumption for each decoding strategy, analyze how hyperparameter tuning impacts energy usage and explain how energy efficiency can be different from decoding latency in the following sections.
4.1 Text Generation Quality vs. Energy Efficiency
(a) Qwen2.5-7B-Instruct
(b) Llama3.1-8B-Instruct
(c) Qwen2.5-14B-Instruct
Figure 1: Relative Standard Deviation of text generation quality and GPU energy usage across decoding strategies.
The text generation quality and energy consumption for various decoding strategies using the Qwen2.5 7B-Instruct model are presented in Table 2. The results were obtained using the best hyperparameter configuration that delivered the highest quality for each decoding method. The corresponding energy consumption for this optimal hyperparameter was averaged across all 5 runs and reported in watt-hours. Similar procedure is used for reporting the results of Llama3.1 8B-Instruct and Qwen2.5 14B-Instruct models in Tables 7 and 8 in Appendix, respectively. To demonstrate the effect of decoding strategies on generative quality and energy consumption, and to test whether the differences between strategies are statistically significant, we used the Friedman[55] test, as our measurements did not follow a normal distribution. This non-parametric test checks for significant differences between two or more groups of repeated measurements. In our case, the repeated measurements are the quality and energy consumption of different decoding strategies in different scenarios (tasks $\times$ models). We applied the Friedman test separately on quality and energy measurements across all models and tasks ($n=15$ blocks, $df=11$). For quality measurements, we obtained $X^{2}=68.77$ with $p=2.09e^{-10}$. For energy measurements, the results were $X^{2}=121.75$ with $p=8.07e^{-21}$. In both cases, the tests were significant, showing that there are clear differences in either quality or energy consumption values of various decoding strategies across the different groups. We also ran the same statistical test for each model and each task separately across different decoding strategies for both quality and energy values. The results are listed in Tables 14 and 15 in Appendix. We observe that the tests were significant in all cases for both quality and energy values, with only one exception in the per-task analysis of quality. This confirms that the choice of decoding strategy consistently has a significant impact on both generative quality and energy consumption, regardless of the specific model or task.
We evaluated the trade-off between quality and energy across different decoding methods using an Efficiency Ratio (ER = Quality/Energy). This metric measures the generative quality achieved per unit of energy consumed, making it suitable for identifying methods that offer high quality while minimizing energy usage. However, care must be taken in interpretation, as the metric can unfairly favor cases where generative accuracy is very low compared to a baseline but energy usage is also low [56].
As shown in Table 2, in both translation tasks, Beam Search and DBS achieve the highest BLEU scores, up to about 42.7 for De→En and 29.9 for En→De, but also draw considerable power, pulling down their overall efficiency ratio. In contrast, AD delivers BLEU scores near—or on par with—those top‐performing beam‐based methods while consuming far less power, thereby attaining the best efficiency ratios. Greedy decoding, along with Top‐p, and Min‐p sampling methods, occupies a middle ground: they offer moderately high BLEU at moderate energy costs, resulting in fairly respectable efficiency.
Similarly, in GSM8K, Beam search achieves the highest exact-match accuracy (0.87), but it requires 14.44 Wh of energy, resulting in a modest efficiency ratio. In contrast, Top-p maintains solid accuracy (0.84) at just 12.78 Wh, yielding one of the highest efficiency ratio among all methods. Several stochastic methods also strike a better balance than beam-based approaches, generating high-quality texts at more moderate energy levels.
@@
Table 2: Text Generation Quality and Average Energy Consumption across Decoding Methods in their best Hyperparameter Setting using Qwen2.5-7B-Instruct with batch size=1. The associated hyperparameters are listed in Table 4. The table uses color coding to highlight key metrics: blue indicates high generative quality, red signifies high energy consumption, and green represents a high efficiency ratio. $\uparrow$ indicates higher is better, $\downarrow$ indicates lower is better.
@@
   De→En $BLEU\uparrow$  41.09 42.71 42.65 22.97 41.03 41.76    41.02 41.26 31.51 31.53 35.68 41.29
  $Wh\downarrow$  9.39 11.13 14.23 22.29 13.26 8.23    9.54 9.37 10.28 10.09 10.05 9.46
  $ER\uparrow$  4.37 3.83 2.99 1.03 3.09 5.07    4.29 4.40 3.06 3.12 3.55 4.36
   En→De $BLEU\uparrow$  26.66 29.88 29.33 19.11 26.94 26.47    26.75 26.96 19.75 18.12 22.13 26.85
  $Wh\downarrow$  13.01 17.87 17.43 21.86 15.40 12.07    13.14 13.24 13.42 13.67 13.54 13.28
  $ER\uparrow$  2.04 1.67 1.68 0.87 1.74 2.19    2.03 2.03 1.47 1.32 1.63 2.02
   GSM8K $Exact-M\uparrow$  0.80 0.87 0.86 0.82 0.81 0.74    0.81 0.84 0.80 0.83 0.84 0.83
  $Wh\downarrow$  13.21 14.44 18.42 17.55 20.28 10.38    13.55 12.78 13.26 12.80 12.98 13.50
  $ER\uparrow$  0.060 0.060 0.046 0.046 0.039 0.071    0.059 0.065 0.060 0.064 0.064 0.061
   HumanEval $pass@1\uparrow$  0.71 0.81 0.79 0.63 0.65 0.69    0.74 0.72 0.54 0.62 0.74 0.78
  $Wh\downarrow$  4.31 4.85 5.96 8.08 4.52 3.16    4.39 3.83 4.87 5.02 4.20 4.03
  $ER\uparrow$  0.164 0.167 0.132 0.077 0.143 0.218    0.168 0.187 0.110 0.123 0.176 0.193
   Wikitext $MAUVE\uparrow$  87.74 91.93 85.22 86.14 92.9 92.32    96.95 94.55 88.87 92.76 90.12 92.53
  $Wh\downarrow$  71.32 73.94 82.87 102.47 80.41 55.26    71.18 71.26 70.49 70.91 70.07 70.93
  $ER\uparrow$  1.23 1.24 1.02 0.84 1.15 1.67    1.36 1.32 1.26 1.30 1.28 1.30
In the HumanEval, Beam Search, DBS, and Min-p tasks, we achieve the highest pass@1 accuracy among all decoding methods. However, Min-p consumes much less energy, resulting in a higher energy efficiency ratio. In Wikitext, we observe that temperature sampling and Top-p sampling achieve the highest MAUVE scores, both exceeding 94, and at the same time rank among the strategies with the highest efficiency ratios, along with AD.
By considering all tasks and calculating the average ranking of decoding strategies based on quality and energy consumption in Qwen2.5 7B-Instruct, AD, Top-p, and Greedy perform best in terms of energy consumption, while Beam Search, Top-p, and DBS perform best in terms of achieved quality. Although AD has the highest efficiency ratio in all cases, in tasks like GSM8K its generative quality is relatively low, which lowers its overall quality ranking. CS, DoLa, and DBS have the highest amount of energy consumption compared to other methods. In contrast, stochastic sampling techniques like Top-p deliver near‐competitive generative quality scores at a fraction of the energy, leading to high efficiency ratios among the tested strategies, especially compared to beam-based deterministic strategies.
@@
To better demonstrate how the choice of decoding strategy affects generative quality and energy consumption, we also employed the Relative Standard Deviation expressed as $\text{RSD}(\%)=\frac{\sigma}{\mu}\times 100$ mathematically. RSD quantifies the variability of the data relative to its mean value. We utilize this statistical metric to provide a normalized measure of variability in text generation quality and energy consumption using various decoding methods across different tasks. Intuitively, a lower RSD indicates greater stability and consistency, while a higher value reflects greater variability. We depicted the RSD of quality and energy consumption of Qwen2.5 7B-Instruct for each task in Figure 1(a) (based on the results from Table 2). It is clear that the selection of decoding strategies significantly affects the variability of energy consumption. Although in some cases, such as GSM8k or WikiText, we do not observe much variation in output quality when using different decoding methods, we do see significant energy deviations in such cases. This highlights the importance of being more energy-conscious when choosing a decoding strategy for our text generation tasks. Choosing an inappropriate strategy may not greatly reduce output quality, but it can lead to a significant increase in energy consumption.
@@
Table 3: Relative Standard Deviation (RSD) and Sharpe Ratio (SR) for each decoding strategy in Qwen2.5-7B-Instruct. RSD is calculated for text generation quality and average energy consumption across different hyperparameter values. SR takes into account the quality and energy consumption of all hyperparameter runs within a decoding method. The table uses color coding to highlight key metrics: blue indicates low RSD in quality, red signifies high RSD in energy consumption, and yellow represents a high SR. $\uparrow$ indicates higher is better.
@@
  De→En BL (RSD%)  0.31 0.46 25.96 9.18 38.36 0.49    18.50 8.96 3.07 2.99 5.81 9.40
  Wh (RSD%)  13.32 11.39 20.08 8.38 35.30 2.25    5.18 2.88 1.22 1.53 2.70 3.09
  $SR\uparrow$  $7.4e^{-3}$ $1e^{-3} 8.8$$8e^{-4}$ $e^{-2}$ $1.3e^{-3}$ $6e^{-2}$  $1.8e^{-2}$ $3.5e^{-2}$ $6.5e^{-2}$ $5.1e^{-2}$ $3.3e^{-2}$ $3.4e^{-2}$
  En→De BL (RSD%)  1.30 2.16 28.86 18.19 51.47 5.41    13.42 13.32 8.04 2.18 4.38 11.62
  Wh (RSD%)  11.09 7.02 16.47 3.96 19.39 4.37    1.49 1.07 0.60 0.92 0.66 2.26
  $SR\uparrow$  $4.5e^{-3}$ $6.8e^{-3}$ $9.4e^{-4}$ $e^{-2}$ $1.2e^{-3}$ $1.3e^{-2}$  $3.3e^{-2}$ $4.3e^{-2}$ $5.3e^{-2}$ $3.5e^{-2}$ $6.5e^{-2}$ $2.1e^{-2}$
  GSM8K EM (RSD%)  3.14 1.77 3.14 5.61 26.67 23.15    3.63 3.99 2.82 2.85 4.12 3.87
  Wh (RSD%)  30.50 20.06 0.31 4.81 17.78 8.61    1.52 0.50 0.90 1.27 1.74 1.18
  $SR\uparrow$  $3.5e^{-5}$ $4.4e^{-5}$ $2.4e^{-3}$ $1.9e^{-4}$ $1.7e^{-5}$ $1.9e^{-4}$  $e^{-3}$ $2.3e^{-3}$ $5e^{-3} 1.1$$1e^{-3}$ $8.8e^{-4}$ $e^{-3}$
  HumanEval ACC (RSD%)  1.79 1.40 11.50 38.86 68.65 9.96    23.94 9.73 2.12 7.00 11.56 12.10
  Wh (RSD%)  25.18 16.30 45.95 11.02 22.35 9.42    19.08 11.59 6.50 11.60 2.58 10.65
  $SR\uparrow$  $1.2e^{-4}$ $1.6e^{-4}$ $2.3e^{-5}$ $2.6e^{-4}$ $5.5e^{-5}$ $6.2e^{-4}$  $1.9e^{-4}$ $3.4e^{-4}$ $4.4e^{-4}$ $2.3e^{-4}$ $1.5e^{-3}$ $4e^{-4}$
  Wikitext MAU (RSD%)  6.56 6.78 11.13 29.00 79.59 3.71    19.97 5.52 7.03 4.45 4.23 6.91
  Wh (RSD%)  8.01 7.17 0.89 1.26 5.15 2.05    0.90 0.87 0.60 1.05 0.80 1.25
  $SR\uparrow$  $4.3e^{-3}$ $3.3e^{-3}$ $2.8e^{-2}$ $2e^{-2}$ $2e^{-3}$ $2.6e^{-2}$  $2.8e^{-2}$ $4.1e^{-2}$ $6.3e^{-2}$ $3.3e^{-2}$ $5.3e^{-2}$ $3.1e^{-2}$
4.1.1 Overall Analysis of Decoding Strategies